# CAST: Corpus-Aware Self-similarity Enhanced Topic modelling

### v1

@@ Abstract
Yanan Ma1thanks:  Corresponding contact email addresses: {yanan.ma, chenhan.yuan, sabine.vanderveer, lamiece.hassan, chenghua.lin, gnenadic}@manchester.ac.uk, chenghao.xiao@durham.ac.uk. Our code will be released upon acceptance.  Chenghao Xiao2 Chenhan Yuan1 Sabine N van der Veer1 Lamiece Hassan1Chenghua Lin1Goran Nenadic1 1The University of Manchester, 2Durham University
@@ 2.1 Candidate Topic Word Embeddings Contextualized on Dataset
Existing neural topic modelling methods like Top2Vec encode document embeddings, taking into account their contextualized meanings. However, when selecting candidate topic words, they encode individual words without their contexts. We posit that, this results in a Contextualization gap, illustrated in Figure 1 (left), i.e., the candidate topic words are not contextualized in any contexts.
@@
In CAST, we first obtain the word embeddings in different documents across the corpus. This involves identifying subword tokens (e.g., those starting with ##) and concatenating them with preceding tokens to form complete words. The word embedding is obtained by averaging the subword embeddings. We then average the word embeddings in different documents to achieve the final word embedding. In this way, the word embedding can preserve the contextual information in the corpus and capture the nuances and variations in word meanings that arise from different contexts. The final word embedding, $E_{\text{final}}$, is given by
$E_{\text{final}}=\frac{1}{P}\sum_{k=1}^{P}\text{Em}([x_{1},x_{2},\ldots,x_{n}]_{k}[V_{w}])$ (1)
where $P$ is the number of different contexts in which the word appears, and Em denotes the last hidden state of the embedding model output for the specified word $w$ in the given sentence. The word embedding $V_{w}$ was calculated by averaging the subword embeddings that form the word.
@@ 2.2 Self-similarity Scores
In contrast to LDA and BERTopic which rely on preprocessing for filtering out functional words, we adopt Top2Vec’s method, which assigns topic candidates to their closest cluster centroids as topic words. Common words, which are equally distant from all clusters, tend to be far from any centroid and are thus discarded. We extend this approach by adding a module to calculate word self-similarity scores across documents, enabling further automatic filtering out of functional words in a post-processing step. This is inspired by the observation of Xiao et al. (2023b) that after contrastive learning, self-similarities of semantic words become higher than a model’s pre-trained checkpoint, while self-similarities of functional words becomes lower. This is referred to as “entourage effect” Xiao et al. (2023a), highlighting that functional word embeddings tend to follow the movement of semantic word embeddings from the same sentence in the embedding space, making the self-similarity for the functional word lower across the documents. This pattern grants us the natural convenience of filtering functional words out - as these are the ones that have lower self-similarities. Given that almost all sentence embedding models used in topic modelling are facilitated by contrastive learning, the pattern becomes universally useful in facilitating our framework.
In CAST, we calculate the words’ cosine similarities with themselves as self-similarity scores. For a word $w$, the self-similarity score $SS_{w}$ is given by
@@
where $E_{w}$ denotes the word embeddings across the corpus.
We then sort the self-similarity scores across the corpus and manually select a threshold to exclude most functional words (those with lower self-similarity scores) while retaining meaningful words as candidate topic words. (e.g. 0.4 in Table 4). More detailed information can be found in Section 5.
@@ 2.3 Document Clustering
To mitigate this issue, we propose to employ Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2018) to reduce the dimensionality of the embeddings and apply HDBSCAN (Campello et al., 2013) to cluster the document embeddings in the reduced dimensions.
@@ 3.2 Baselines
BERTopic employs a hybrid approach that combines embedding models and a class-based tf-idf method for topic words assignment. Similar to Top2Vec, it encodes and clusters the document embeddings. For each cluster, however, it diverges by using tf-idf to identify topic words from documents within the same clusters.
@@
In our experiment, we loaded the default LDA from the OCTIS library and used the default parameters of TopClus. We employed two sentence-transformers, all-MiniLM-L6-v2 and all-mpnet-base-v2, to run the Top2Vec, BERTopic, and CAST. We reduced the embedding dimensionality to five and adjusted the min_cluster_size and the #Topics to detect based on the dataset size (see Table 6), with larger datasets having a larger min_cluster_size value and more topics to detect. We kept and fixed the other parameters for UMAP and HDBSCAN across all models for a fair comparison (see Appendix C). For each model, all the results were averaged across five independent runs across three different numbers of topics in the datasets (15 runs per model in total) by using one NVIDIA A100 GPU.
@@
Table 1: We evaluated all models using traditional topic coherence (TC) and topic diversity (TD), alongside large language model-based metrics: LLM-TC and LLM-TD. For each model, all the results were averaged across five independent runs across three different numbers of topics in the datasets (15 runs per model in total). The best two values under each metric were highlighted in green with darker colors indicating greater value. means the model does not use the self-similarity module to filter words.
@@ 3.3 Evaluation Metrics
Qualitative evaluation. We further examined the ability of topic models to generate meaningful words over noisy words as a metric to evaluate the models. To achieve this, we randomly selected five labels and manually matched the most relevant topics generated by all models. We then manually reviewed the topic words, underlining and italicizing those irrelevant to the topic, which we identified as noisy words (see Table 2). We only conducted the qualitative evaluation on 20NewsGroups and BBC News datasets, as they are human-labelled datasets with topic labels.
@@ 4 Results
In terms of TC, CAST models exhibit consistently high TC scores on both traditional NPMI and LLM-based coherence metrics across all datasets. Among the two variants, the MPNET variant achieves the highest NPMI and LLM-TC score on the 20NewsGroups and Elon Musk dataset. The MiniLM variant demonstrates consistent and competitive coherence on all datasets, with the best LLM-TC score on BBC News, and the second-best on 20NewsGroups.
@@
The vanilla CAST models (marked with *), without using the self-similarity module, generally outperform baseline models, underscoring the efficacy of the contextualized embedding module to improve the model performance. However, these vanilla models exhibit marginally lower coherence scores while maintaining robust diversity metrics compared with the full CAST models. This suggests that the self-similarity module, by filtering functional words, could enhance coherence at a slight cost to topic word diversity. We further explain this in Section 5.
Qualitative results. Table 2 further analyzes models’ ability to generate meaningful topic words over noisy words. In general, Top2Vec and our proposed model demonstrate a consistent ability to generate coherent and contextually relevant topic words. However, Top2Vec fails to identify Topic 4 (Space) in the 20NewsGroups dataset. This is potentially due to Top2Vec’s approach of merging the smallest cluster with the most similar ones to achieve the desired number of topics. While this method can effectively merge similar topics, it could also lead to pollution to larger clusters and potentially ignore certain topics. In contrast, CAST selects the top-$n$ clusters according to their sizes to determine the desired number of topics. This approach ensures that significant topics are preserved and minimizes the risk of smaller clusters distorting the embeddings of larger, meaningful clusters.
In terms of LDA and BERTopic, they frequently generate noisy and semantically irrelevant words potentially because they do not leverage the semantic richness offered by word embeddings to identify topic words. TopClus and BERTopic models exhibit significant gaps in topic coverage. Specifically, BERTopic struggles to distinguish between Topics 2 and 4 in the 20NewsGroups dataset, leading to a conflation of topic words. This indicates a limitation in the model’s ability to maintain distinct topic boundaries, which is crucial for generating precise and meaningful topic representations. Apart from the topics shown in the table, we found that TopClus tend to cluster meaningless adjectives and verbs together for example ‘good’, ‘long’, ‘big’, ‘small’, ‘large’. More detailed examples can be found in Appendix D.
@@
Table 3: We tested the models’ performance in dealing with noisy data on the unprocessed 20 Newsgroups dataset fetched from sklearn datasets, with the number of topics set to 10. Results were averaged across five independent runs for each model with the best value in each metric highlighted in bold and the second best underlined. We manually chose one typical topic with five topic words to better show the models’ ability to deal with noisy data.
Figure 2: Ablation analysis of varying self-similarity thresholds on the performance of our MPNET variant model on the 20NewsGroups (nr_topics = 10) and BBC News (nr_topics = 5) datasets. All results were averaged across five independent runs for each threshold value. The analysis was constrained to a maximum threshold of 0.7 because the model could not identify a sufficient number of topic words after this threshold.
@@ 4.1 Dealing with Noisy Data
In real-world settings, data are full of noise. Therefore, we evaluated the performance of the models in handling noisy data in this section. Table 3 illustrates the results of our MPNET variant model and corresponding baseline models tested on the 20 Newsgroups dataset [^3] https://scikit-learn.org/0.15/datasets/twenty_newsgroups.html as an example, with punctuations and stopwords retained.
The results demonstrate that LDA and BERTopic struggle to effectively handle noisy data due to their limitations for topic word assignment. Similarly, TopClus, despite its ability to partially mitigate this issue, tends to cluster functional words, such as verbs and adjectives, as topic words, which may not accurately capture the underlying themes. In contrast, Top2Vec and CAST exhibit robustness in extracting meaningful topic words from noisy data.
Notably, while BERTopic performed second-best on NPMI, its LLM-TC score is quite low, as evident from the topic word examples. This shows that meaningless topics comprising several stopwords may sometimes receive higher NPMI scores due to the common co-occurrence of stopwords like ‘the”, “to”, “and”, “is” within documents. This observation aligns with previous criticisms that word co-occurrence-based metrics, such as NPMI, are incompatible with recently developed neural topic models (Hoyle et al., 2021; Doogan and Buntine, 2021). Therefore, it is crucial to use suitable coherence metrics to evaluate recently developed neural topic models, where a potential solution is using LLMs to evaluate the models, as shown in the paper.
@@
Table 4: Example of words and their self-similarity (SS) scores from the 20NewsGroups dataset. We display the top 10 words with the highest SS scores, alongside the top 10 words with SS scores below three thresholds, namely, 0.5, 0.4, and 0.3.
@@ 5 Ablation Analysis
In this section, we ablate the self-similarity threshold to evaluate how varying self-similarity affects TC and TD. We tested our MPNET variant model on the 20NewsGroups and BBC News datasets.
As shown in Figure 2, increasing the self-similarity thresholds initially enhances the model’s performance, but experiences a marked decline after a higher threshold. As tested on the 20NewsGroups and BBC News datasets, TC and LLM-TC remain competitive and peak around a threshold of 0.4 in most cases, while TD and LLM-TD remain stable with minor fluctuations up to this point. However, beyond a self-similarity score of 0.5, the coherence and diversity metrics experience a significant decline. This pattern indicates that an optimal range of self-similarity scores can increase topic coherence and diversity, but excessively high thresholds may filter out meaningful words, thereby reducing the performance.
@@
Additionally, when setting self-similarity to 0, ablating the use of the self-similarity module, CAST still outperforms other models under the same conditions in most cases (see Table 1). This demonstrates the efficacy of the contextual embedding method employed by CAST in enhancing the identification of coherent and diverse topics.
@@ 6 Related Work
Neural Topic models. Recent neural topic models such as Top2Vec and BERTopic combine embedding techniques with clustering algorithms. They have demonstrated superior performance in identifying coherent and diverse topics (Angelov, 2020; Grootendorst, 2022; Egger and Yu, 2022), and have been widely used in various domains (Ghasiya and Okamura, 2021; Jeon et al., 2023; Wang et al., 2023).
Our model uses similar methods as Top2Vec and BERTopic for dimensionality reduction and embedding clustering. However, it differs in attaining the candidate cluster topic word embeddings by averaging the embeddings of the same word in different contexts across the corpus, to capture contextual information of the corpus. In contrast, Top2Vec typically extracts static word embeddings from pre-trained embeddings. BERTopic, instead of building word embeddings and finding candidate words whose embeddings are close to the topic embeddings, identifies topic words through a clustering-based tf-idf approach. Furthermore, our method introduces self-similarity scores as a novel mechanism for filtering candidate topic words. Typically, functional words exhibit lower self-similarity scores in their embeddings compared to semantically meaningful words. By establishing an appropriate threshold, we can effectively exclude most functional words while retaining semantically significant terms. This could enhance topic coherence and improve interpretability for end-users.
Contrastive Learning. In the contextualized language model era, contrastive learning has been facilitated by sentence embedding models used for topic modelling Oord et al. (2018); Gao et al. (2021); Reimers and Gurevych (2019). Our method is largely facilitated by properties brought by contrastive learning, including the different self-similarity patterns for functional tokens and semantic tokens; and the usage of contextualized embeddings to serve as candidates for centroid words Ethayarajh (2019); Xiao et al. (2023b). This is largely inspired by the “entourage effect” found in Xiao et al. (2023b, a), highlighting that words in the same sentence all follow semantic tokens, wherever they travel, motivating the usage of cluster-specific candidate contextualized embeddings in our work. To the best of our knowledge, previous work in topic modelling has not been fully leveraged these useful properties inherent in sentence embeddings brought by contrastive learning.

### v2

@@ Abstract
Yanan Ma1thanks:  Corresponding contact email addresses: {yanan.ma, chenhan.yuan, sabine.vanderveer, lamiece.hassan, chenghua.lin, gnenadic}@manchester.ac.uk, chenghao.xiao@durham.ac.uk. Our code is available at: https://github.com/yananma1029/CAST.  Chenghao Xiao2 Chenhan Yuan1 Sabine N van der Veer1 Lamiece Hassan1Chenghua Lin1Goran Nenadic1 1The University of Manchester, 2Durham University
@@ 2.1 Candidate Topic Word Embeddings Contextualized on Dataset
Existing neural topic modelling methods like Top2Vec encode document embeddings, taking into account their contextualized meanings. However, when selecting candidate topic words, they encode individual words without their contexts. We posit that this results in a Contextualization gap, illustrated in Figure 1 (left), i.e., the candidate topic words are not contextualized in any context.
@@
In CAST, we first extract the word embeddings from different documents across the corpus and then average them to obtain the final word embedding. In this way, the word embedding can preserve the contextual information in the corpus and capture the nuances and variations in word meanings that arise from different contexts.
The final word embedding, $E_{\text{final}}$, is defined as the average of the contextualized word embeddings of a word $w$ across all sequences in a corpus. Formally:
$E_{\text{final}}=\frac{1}{P}\sum_{i=1}^{P}C_{wi}$ (1)
where $C_{wi}$ denotes the contextualized word embedding of the word $w$ in the $i$-th sequence. $P$ is the number of sequences that contain $w$.
Given a sentence $S$, the following steps are performed to compute $C_{wi}$:
  • 1.
    Tokenize $S$ into subwords $\{t_{1},t_{2},\dots,t_{n}\}$.
  • 2.
    For each word $w$, composed of subwords $\{t_{i_{1}},t_{i_{2}},\dots,t_{i_{k}}\}$ (e.g. ‘‘happi’’, ‘‘##ness’’), its $C_{wi}$ is calculated as:
    $C_{wi}=\frac{mean\left(\mathbf{e}_{i_{1}},\mathbf{e}_{i_{2}},\ldots,\mathbf{e}_{i_{k}}\right)}{\left|mean\left(\mathbf{e}_{i_{1}},\mathbf{e}_{i_{2}},\ldots,\mathbf{e}_{i_{k}}\right)\right|}$
where $mean$ denotes the element-wise mean of the subword embeddings, and ||$\cdot$|| ⋅ | represents the $L2$ norm.
@@ 2.2 Self-similarity Scores
In contrast to LDA and BERTopic which rely on preprocessing for filtering out functional words, we adopt Top2Vec’s method, which assigns topic candidates to their closest cluster centroids as topic words. Common words, which are equally distant from all clusters, tend to be far from any centroid and are thus discarded. We extend this approach by adding a module to calculate word self-similarity scores across documents, enabling further automatic filtering out of functional words in a post-processing step. This is inspired by the observation of Xiao et al. (2023b) that after contrastive learning, self-similarities of semantic words become higher than a model’s pre-trained checkpoint, while self-similarities of functional words become lower. This is referred to as “entourage effect” Xiao et al. (2023a), highlighting that functional word embeddings tend to follow the movement of semantic word embeddings from the same sentence in the embedding space, making the self-similarity for the functional word lower across the documents. This pattern grants us the natural convenience of filtering functional words out - as these are the ones that have lower self-similarities. Given that almost all sentence embedding models used in topic modelling are facilitated by contrastive learning, the pattern becomes universally useful in facilitating our framework.
In CAST, we calculate each word’ cosine-similarity with itself as its self-similarity score. For a word $w$, its self-similarity score $SS_{w}$ is given by
@@
where $E_{w}=\{C_{w1},C_{w2},\dots,C_{wi}\}$ is a list of contextualized word embeddings of word $w$ across the corpus.
We then sort the self-similarity scores and manually select a threshold to exclude most functional words (those with lower self-similarity scores) while retaining meaningful words as candidate topic words. For example, setting threshold = 0.3 could exclude functional words such as “<pad>” and “due” (see Table 4).
@@ 2.3 Document Clustering
To mitigate this issue, we employed Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2018) to reduce the dimensionality of the embeddings and applied HDBSCAN (Campello et al., 2013) to cluster the document embeddings in the reduced dimensions.
@@ 3.2 Baselines
BERTopic employs a hybrid approach that combines embedding models and a class-based tf-idf method for topic word assignment. Similar to Top2Vec, it encodes and clusters the document embeddings. For each cluster, however, it diverges by using tf-idf to identify topic words from documents within the same clusters.
@@
In our experiment, we loaded the default LDA from the OCTIS library and used the default parameters of TopClus. We employed two sentence-transformers, all-MiniLM-L6-v2 and all-mpnet-base-v2, to run Top2Vec, BERTopic, and CAST. We reduced the embedding dimensionality to five and adjusted the min_cluster_size and the #Topics based on the dataset size (see Table 6), with larger datasets having a larger min_cluster_size value and more topics to detect. We kept and fixed the other parameters for UMAP and HDBSCAN across all models for a fair comparison. For each model, all the results were averaged across five independent runs across three different numbers of topics in the datasets (15 runs per model in total) by using one NVIDIA A100 GPU. See Appendix C for a more detailed implementation description.
@@
Table 1: We evaluated all models using traditional topic coherence (TC) and topic diversity (TD), alongside large language model-based metrics: LLM-TC and LLM-TD. For each model, all the results were averaged across five independent runs across three different numbers of topics in the datasets (15 runs per model in total). The best two values under each metric were highlighted in green with darker colors indicating greater value. means the vanilla model that does not use the self-similarity module to filter words.
@@ 3.3 Evaluation Metrics
Qualitative evaluation. We further examined the ability of topic models to generate meaningful words over noisy words as a metric to evaluate the models. To achieve this, we randomly selected five labels and manually matched the most relevant topics generated by all models. We then manually reviewed the topic words, underlining and italicizing those irrelevant to the topic, which we identified as noisy words (see Table 2). We only conducted the qualitative evaluation on 20NewsGroups and BBC News datasets, as they are human-labelled datasets with topic labels (e.g. sports, politics).
@@ 4 Results
In terms of TC, CAST models exhibit consistently high TC scores on both traditional NPMI and LLM-based coherence metrics across all datasets. Among the two variants, CAST-MPNET achieves the highest NPMI and LLM-TC scores on the 20NewsGroups and Elon Musk datasets. CAST-MiniLM demonstrates consistent and competitive coherence on all datasets, with the best LLM-TC score on BBC News, and the second-best on 20NewsGroups.
@@
The vanilla CAST models (marked with *), which ablate the self-similarity module, still outperform baseline models in most cases, underscoring the efficacy of the contextualized embedding module to improve the model performance. However, compared with the full CAST models, the vanilla models exhibit marginally lower coherence scores while maintaining robust diversity metrics. This suggests that the self-similarity module, by filtering out functional words, could enhance coherence at a slight cost to topic word diversity. We further explain this in Section 5.
Qualitative results. Table 2 further analyzes models’ ability to generate meaningful topic words over irrelevant words. In general, Top2Vec and CAST are consistent in generating coherent and contextually relevant topic words. However, Top2Vec fails to identify Topic 4 (Space) on 20NewsGroups. Furthermore, Table 8 reveals that Top2Vec-MPNET occasionally generates less meaningful topic words, such as “inform”, “topic”, and “implement” in the first two topics, and the second topic is too ambiguous, making it hard to distinguish between medicine and business. This is potentially due to Top2Vec’s approach of merging the smallest cluster with the most similar ones to achieve the desired number of topics. While this method can effectively merge similar topics, it could also lead to pollution to larger clusters and potentially ignore certain topics. In contrast, CAST selects the top-$n$ clusters according to their sizes to determine the desired number of topics. This approach ensures that significant topics are preserved and minimizes the risk of smaller clusters distorting the embeddings of larger, meaningful clusters.
In terms of LDA and BERTopic, they frequently generate noisy and semantically irrelevant words despite their limited meaningfulness. This is potentially because these two models do not leverage the semantic richness offered by word embeddings to identify topic words. TopClus and BERTopic models exhibit significant gaps in topic coverage. Specifically, BERTopic struggles to distinguish between Topics 2 and 4 in the 20NewsGroups dataset, leading to a conflation of topic words. This indicates a limitation in the model’s ability to maintain distinct topic boundaries. TopClus tends to cluster meaningless adjectives and verbs together as a topic. For example, “good”, “long”, “big”, “small”, “large” in Table 8.
@@
Table 3: We tested the models’ performance in dealing with noisy data on the non-preprocessed 20 Newsgroups dataset fetched from sklearn, with the number of topics set to 10. Results were averaged across five independent runs for each model with the best value in each metric highlighted in bold and the second best underlined. We manually chose one typical topic with five topic words to better show the models’ ability to deal with noisy data.
Figure 2: Ablation analysis of varying self-similarity thresholds on CAST-MPNET for the 20NewsGroups (nr_topics = 10) and BBC News (nr_topics = 5) datasets. All results were averaged across five independent runs for each threshold value. The analysis was constrained to a maximum threshold of 0.7 because the model could not identify sufficient topic words after this threshold.
@@ 4.1 Dealing with Noisy Data
In real-world settings, data are full of noise. Therefore, we evaluated the performance of the models in handling noisy data in this section. Table 3 illustrates the results of CAST-MPNET and its corresponding baseline models tested on the non-preprocessed 20 Newsgroups dataset [^3] https://scikit-learn.org/0.15/datasets/twenty_newsgroups.html as an example.
The results demonstrate that LDA and BERTopic struggle to effectively handle noisy data due to their limitations for topic word assignment. Similarly, TopClus, despite its ability to partially mitigate this issue, tends to cluster functional words, such as verbs and adjectives, as topic words, which may not accurately capture the underlying topics. In contrast, Top2Vec and CAST exhibit robustness in extracting meaningful topic words from noisy data. Both methods do not need preprocessing to identify topic words, as common words—equidistant from all clusters—are typically far from any centroid and are thus discarded (see Section 2.2). However, some less common function words such as “inform” and “implement” may still be wrongly selected by Top2Vec (see Table 8). CAST extends Top2Vec with an automatic post-processing self-similarity module, which further removes less meaningful function words. For example, the self-similarity scores for “inform” and “implement” are 0.218 and 0.26 respectively, and thus are excluded when setting the self-similarity threshold to be 0.4. We further explain the impact of varying thresholds on CAST’s performance in Section 5.
On the other hand, while BERTopic performed second-best on NPMI, its LLM-TC score is quite low, as evident from the topic word examples. This shows that meaningless topics comprising several stopwords may sometimes receive higher NPMI scores due to the common co-occurrence of stopwords like “the”, “to”, “and”, “is” within documents. This observation aligns with previous criticisms that word co-occurrence-based metrics, such as NPMI, are incompatible with recently developed neural topic models (Hoyle et al., 2021; Doogan and Buntine, 2021). Therefore, it is crucial to use suitable coherence metrics to evaluate recently developed neural topic models, where a potential way is to use LLMs to evaluate the models, as shown in the paper.
@@
Table 4: Example of words and their self-similarity (SS) scores from the 20NewsGroups dataset. We present the top 10 words with the highest SS scores, along with the top 10 words whose SS scores fall below three thresholds: 0.5, 0.4, and 0.3.
@@ 5 Ablation Analysis
In this section, we ablate the self-similarity thresholds to evaluate how varying thresholds affect TC and TD. We tested CAST-MPNET on the 20NewsGroups and BBC News datasets.
As shown in Figure 2, increasing the self-similarity threshold initially enhances the model’s performance, but experiences a marked decline after a higher threshold. As tested on the 20NewsGroups and BBC News datasets, TC and LLM-TC remain competitive and peak around a threshold of 0.4 in most cases, while TD and LLM-TD remain stable with minor fluctuations up to this point. However, beyond a self-similarity score of 0.5, the coherence and diversity metrics experience a significant decline. This pattern indicates that an optimal range of self-similarity scores can increase topic coherence and diversity, but excessively high thresholds may filter out meaningful words, thereby reducing the performance.
@@
Additionally, when setting self-similarity to 0, ablating the use of the self-similarity module, CAST still outperforms other models under the same conditions in most cases (see the vanilla models in Table 1). This demonstrates the efficacy of the contextual embedding method employed by CAST in enhancing the identification of coherent and diverse topics.
@@ 6 Related Work
Neural Topic models. Recent neural topic models such as Top2Vec and BERTopic combine embedding techniques with clustering algorithms. They have demonstrated superior performance in identifying coherent and diverse topics (Angelov, 2020; Grootendorst, 2022; Egger and Yu, 2022), and have been widely used in various domains (Ghasiya and Okamura, 2021; Jeon et al., 2023; Wang et al., 2023b).
Our approach uses similar methods as Top2Vec and BERTopic for dimensionality reduction and embedding clustering. However, it differs in attaining the candidate topic word embeddings by averaging the embeddings of the word in different contexts across the corpus, to capture contextual information of the corpus. In contrast, Top2Vec typically extracts static word embeddings from pre-trained embeddings. BERTopic, instead of building word embeddings and finding candidate words whose embeddings are close to the topic embeddings, identifies topic words through a clustering-based tf-idf approach. Furthermore, our method introduces self-similarity scores as a novel mechanism for filtering candidate topic words. Typically, functional words exhibit lower self-similarity scores in their embeddings compared to semantically meaningful words. By establishing an appropriate threshold, we can effectively exclude most functional words while retaining semantically significant terms. This could enhance topic coherence and improve interpretability for end-users.
LLM-based topic models. Recently LLM-based topic models (Mu et al., 2024; Wang et al., 2023a), typically relying on prompt-based strategies to guide LLMs to identify latent topics, have shown strong potential for generating coherent topics. Our approach is complementary to the LLM-based methods by offering a lightweight and cost-effective alternative. By leveraging pre-trained embeddings and contextualization techniques, it can produce coherent and interpretable results without the extensive computational cost required by LLMs, making it more accessible for scenarios with limited resources.
Contrastive Learning. In the contextualized language model era, contrastive learning has been facilitated by sentence embedding models used for topic modelling Oord et al. (2018); Gao et al. (2021); Reimers and Gurevych (2019). Our method is largely facilitated by properties brought by contrastive learning, including the different self-similarity patterns for functional tokens and semantic tokens; and the usage of contextualized embeddings to serve as candidates for centroid words Ethayarajh (2019); Xiao et al. (2023b). This is largely inspired by the “entourage effect” found in Xiao et al. (2023b, a), highlighting that words in the same sentence all follow semantic tokens, wherever they travel, motivating the usage of cluster-specific candidate contextualized embeddings in our work. To the best of our knowledge, previous work in topic modelling has not fully leveraged these useful properties inherent in sentence embeddings brought by contrastive learning.